Find four consecutive even integers such that if the sum of the first and third is multiplied by 6 the result is 8 more than 10 times the 4th.
can you please show me the formula also because i have looked all through my text book and it doesnt explain it in detail and skips steps.

Answer by Earlsdon(6103)

Let the first even integer be x.
The second consecutive even integer will be (x + 2)
The third one will be (x + 4)
And the fourth one will be (x + 6)
Now write the necessary equations.
The sum of the 1st and 3rd multiplied by 6 is:
6[x + (x + 4)]
The result is (=): 8 more (+8) than 10 times the fourth 10(x + 6) or:
10(x + 6) + 8
Put it all together and solve for x:
6[x + (x + 4)] = 10(x + 6) + 8 Simplify.
6[2x + 4] = 10x + 60 + 8
12x + 24 = 10x + 68 Subtract 10x from both sides.
2x + 24 = 68 Subtract 24 from both sides.
2x = 44 Divide both sides by 2.
x = 22 This is the first even integer.
x + 2 = 24 This is the second consecutive even integer.
x + 4 = 26 This is the third consecutive even integer.
x + 6 = 28 This is the fourth consecutive even integer.
Check:
6[x + (x + 4)] = 6[22 + (22 + 4)] = 6[22 + 26] = 6[48] = 288
10(x + 6) + 8 = 10(22 + 6) + 8 = 10(28) + 8 = 280 + 8 = 288